### v1

# lemmasection

Abstract
ECG foundation models are increasingly popular due to their adaptability across various tasks. However, their clinical applicability is often limited by performance gaps compared to task-specific models, even after pre-training on large ECG datasets and fine-tuning on target data. This limitation is likely due to the lack of an effective post-training strategy. In this paper, we propose a simple yet effective post-training approach to enhance ECGFounder, a state-of-the-art ECG foundation model pre-trained on over 7 million ECG recordings. Experiments on the PTB-XL benchmark show that our approach improves the baseline fine-tuning strategy by 1.2%–3.3% in macro AUROC and 5.3%–20.9% in macro AUPRC. Additionally, our method outperforms several recent state-of-the-art approaches, including task-specific and advanced architectures. Further evaluation reveals that our method is more stable and sample-efficient compared to the baseline, achieving a 9.1% improvement in macro AUROC and a 34.9% improvement in macro AUPRC using just 10% of the training data. Ablation studies identify key components, such as stochastic depth and preview linear probing, that contribute to the enhanced performance. These findings underscore the potential of post-training strategies to improve ECG foundation models, and we hope this work will contribute to the continued development of foundation models in the ECG domain.
##footnotetext: These authors contributed equally.**footnotetext: Corresponding authors: Wei Zhao (zhaowei@nhc.gov.cn), Xiaohan Fan (fanxiaohan@fuwaihospital.org), Ya Zhou (zhouya@fuwai.com) footnotetext: Supervision: Wei Zhao
Keywords: Electrocardiogram, Post-training strategy, ECGFounder, Model adaptation, Foundation models
1 Introduction
Electrocardiography (ECG) is a widely used, non-invasive technique for the screening, diagnosis, monitoring, and risk prediction of cardiovascular diseases. ECG captures the heart’s electrophysiological activity, offering valuable insights into cardiac function and health. In recent years, the rapid evolution of artificial intelligence, particularly deep learning, has revolutionized ECG analysis. Deep neural networks (DNNs) have demonstrated expert-level accuracy in detecting a wide range of cardiac abnormalities, such as arrhythmias (Hannun et al., 2019; Ribeiro et al., 2020). These models are increasingly deployed in clinical settings to assist cardiologists, alleviate diagnostic workloads, and enhance decision-making (Jiang et al., 2024). Furthermore, DNNs have demonstrated the ability to identify subtle signal patterns undetectable by human experts (Somani et al., 2021; Siontis et al., 2021), enabling the diagnosis of conditions like structural heart disease (Poterucha et al., 2025), anemia (Kwon et al., 2020), and diabetes (Kulkarni et al., 2023). This has expanded the diagnostic potential of ECG far beyond its traditional role.
Recently, foundation models have emerged as a transformative approach for building generalizable and transferable models across diverse clinical tasks (Moor et al., 2023). Trained on large-scale datasets, these models can be fine-tuned for various downstream tasks (Bommasani et al., 2021). In the ECG domain, foundation models such as ECG-FM (McKeen et al., 2024), HuBERT-ECG (Coppola et al., 2024), and ECGFounder (Li et al., 2025) have shown broad applicability across ECG-related tasks. The recently released foundation model, ECGFounder, built on the Harvard-Emory ECG Database (Koscova et al., 2024), which includes over 10 million clinically annotated ECG recordings covering 150 types of cardiac abnormalities, has demonstrated strong performance and generalizability across multiple applications.
Despite significant advancements in ECG foundation models, substantial performance gaps persist when fine-tuned for specific diagnostic tasks, especially compared to recent task-specific models. For example, a recent study (Coppola et al., 2024) reveals that HuBERT-ECG BASE, while performing well across several tasks on the PTB-XL benchmark, falls short when compared to those specialized models (Strodthoff et al., 2021). Encouraged by the open-source efforts of Li et al. (2025), we conducted a series of experiments using the PTB-XL dataset. However, we observed that the fine-tuning performance of even the state-of-the-art foundation model ECGFounder remains suboptimal. This highlights a key issue: when ECG foundation models pre-trained on one large dataset are applied to ECG signals from other data sources, they often fail to match the performance of specialized models. Moreover, compared to recent advanced architectures—such as multi-scale models (Shi et al., 2023) and state-space models (Zhang et al., 2023; Behrouz et al., 2024)—the performance of these foundation models lags significantly, raising concerns about their practical applicability in real-world clinical settings.
To address these limitations, we propose a simple yet effective post-training strategy based on the recently released ECGFounder model. We use the PTB-XL all-label classification benchmark as a performance example. As illustrated in Figure 1, while pre-training on over 7 million ECG recordings brings substantial improvements over training from scratch, the benefits of pre-training are closely matched by our post-training approach, even when applied to randomly initialized models. When applied to the pre-trained ECGFounder, our proposed post-training strategy leads to remarkable performance gains of 3.2% in AUROC and 14.3% in AUPRC (Figures 1(a) and 1(b)). Furthermore, our experiments show that, compared to state-of-the-art task-specific models, our method achieves superior classification performance on several PTB-XL benchmark tasks. These results underscore the complementary advantages of the proposed post-training strategy, demonstrating its ability to address the performance gaps observed in ECG foundation models.
Figure 1: Performance comparison on PTB-XL 71-label classification benchmark. (a) Marco AUROC results comparing different training strategies. (b) Macro AUPRC results under the same experimental setup. “Random Initialization” denotes training from scratch, while “Pre-training” utilizes transferred weights. Both subfigures contrast baseline fine-tuning methods (left) with our proposed post-training approach (right).
The proposed post-training strategy is grounded in two key insights. First, ECG signals inherently exhibit information redundancy, where neighboring time points and heartbeat cycles are often predictable from each other (Zhou et al., 2025). To address this, we employ stochastic depth (Huang et al., 2016) to reduce redundancy and enhance the model’s robustness. Second, pre-training provides a valuable initialization that generally improves model performance. However, while the early layers of the model retain the pre-trained weights, the final classification head is typically initialized randomly. To optimize this, we introduce preview linear probing during post-training to better initialize the classification head, thereby boosting task-specific performance. Although these two components—stochastic depth and preview linear probing—are relatively simple, they play a crucial role in driving improvements in the performance of the ECG foundation model.
To explore the effectiveness of the proposed post-training strategy, we conducted a series of experiments, including varying sample sizes and training length, and an ablation study. By evaluating performance under diverse conditions, we found that the strategy is both stable and sample-efficient, showing its capacity to enhance model performance not only with full datasets but also in scenarios with limited labeled data. These experiments provide valuable insights, demonstrating the promising performance and potential applicability of the method.
In summary, the main contributions of this work are as follows:
  • 1.
    Post-Training Strategy for ECGFounder. We propose a simple yet effective post-training strategy to enhance ECGFounder, a recently released state-of-the-art ECG foundation model. This strategy provides a crucial complement to the original framework, facilitating more efficient adaptation to diverse downstream ECG tasks.
  • 2.
    Significant Performance Gains. Experiments on the PTB-XL benchmark demonstrate that our method outperforms the baseline fine-tuning strategy from the original work across multiple tasks, by 1.2%–3.3% in macro AUROC and 5.3%–20.9% in macro AUPRC. Furthermore, we explore the properties of the proposed method under varying sample sizes and training epochs. Notably, with only 10% of the training data, the proposed method achieves remarkable improvements of 9.1% in macro AUROC and 34.9% in macro AUPRC, highlighting its effectiveness in data-scarce scenarios.
  • 3.
    Superior Performance. Our method surpasses several recent state-of-the-art task-specific and advanced models across multiple ECG classification tasks, highlighting the potential of ECG foundation models when equipped with an effective post-training strategy.
  • 4.
    Practical Implications. Recent studies have reported sub-optimal performance of fine-tuned ECG foundation models on the PTB-XL benchmark, raising concerns regarding their clinical applicability. Our results suggest that this limitation likely stems from the absence of an effective post-training strategy, which could address the gap and enable real-world clinical adoption.
2 Related Work
2.1 ECG Classification and Evaluation Benchmarks
ECG classification is one of the most important tasks in ECG analysis, crucial for many critical applications, including the detection of arrhythmia (Hannun et al., 2019; Ribeiro et al., 2020; Hughes et al., 2021), structural heart disease (Poterucha et al., 2025), anemia (Kwon et al., 2020), and diabetes (Kulkarni et al., 2023). Due to their excellent performance in both accuracy and diagnostic speed (Jiang et al., 2024), deep learning methods have gained significant popularity in these tasks. Successful model architectures in these areas range from convolutional neural networks (CNNs) and Transformer to state-space models and ensemble approaches (Strodthoff et al., 2021; Cheng et al., 2023; Shi et al., 2023; Zhang et al., 2023; Behrouz et al., 2024). Recent advanced architectures have proven effective in modeling the time-series nature of ECG signals, including multi-scale models like MSW-Transformer (Cheng et al., 2023) and MULTIRESNET (Shi et al., 2023), as well as advanced state-space models like SPACETIME (Zhang et al., 2023) and Chimera (Behrouz et al., 2024).
However, comparing the performance of various ECG classification methods presents several challenges. While numerous large ECG datasets have been collected, many remain private (Somani et al., 2021; Siontis et al., 2021). Recent efforts to address this limitation have led to the release of several large, publicly available ECG datasets (Clifford, 2024). Nonetheless, a fair comparison requires more than just raw datasets. Clearly defined usage guidelines, such as pre-defined train-test-splits, label sets, and target metrics, are essential to ensure valid comparisons (Strodthoff, 2024). Among these, the PTB-XL dataset (Wagner et al., 2020) stands out as one of the most comprehensive benchmarks, providing essential data for evaluating various ECG classification methods. Furthermore, based on this dataset, Strodthoff et al. (2021) presents benchmark results for several methods (Ismail Fawaz et al., 2020; Wang et al., 2017), providing an important baseline for comparing different ECG classification algorithms.
2.2 ECG Foundation Models
Recent advances in foundation models for natural language processing have inspired similar approaches for health data (Clifford, 2024). In the ECG domain, several studies have explored building foundation models through large-scale pre-training followed by evaluation across diverse downstream tasks. Representative examples include ECG-FM (McKeen et al., 2024), HuBERT-ECG (Coppola et al., 2024), and ECGFounder (Li et al., 2025). Specifically, ECG-FM is a Transformer-based model pre-trained on over 1 million 12-lead ECGs from multiple sources; HuBERT-ECG is another Transformer-based model pre-trained on more than 9 million 12-lead ECGs covering 164 cardiovascular conditions; and ECGFounder is a CNN-based model pre-trained on over 7 million 12-lead ECGs annotated with 150 diagnostic categories. Importantly, the codes and pre-trained model weights of all three models have been publicly released.
Comparing the performance of different ECG foundation models would provide valuable insights for practitioners. However, conducting a fair comparison is challenging, especially for ECG foundation models due to their higher computational cost. As discussed in Section 2.1, the PTB-XL benchmark offers a fair platform for evaluating various algorithms. Nonetheless, the pre-training dataset for ECG-FM uses random splits of the PhysioNet 2021 dataset (Reyna et al., 2021; Goldberger et al., 2000), which includes the PTB-XL dataset, potentially limiting its applicability for evaluation on the PTB-XL benchmark. For HuBERT-ECG, the authors have reported performance results on PTB-XL, enabling direct comparison. For ECGFounder, we follow the recommended fine-tuning strategy (Li et al., 2025) based on the released model weights to evaluate performance on the PTB-XL benchmark.
2.3 Regularization in Deep Learning
Deep learning models often require significant capacity to capture complex representations, making them susceptible to overfitting, especially when the available training data is limited. Regularization is a widely used technique to mitigate overfitting and improve model generalization. In general, regularization refers to any technique applied in the model architecture, learning process, or inference to alleviate the limitations of insufficient training data (Moradi et al., 2020). A common regularization method is weight decay (Hanson and Pratt, 1988), which penalizes large weights during each update step. Loshchilov and Hutter (2019) demonstrated that weight decay, combined with a cosine annealing learning rate schedule (Loshchilov and Hutter, 2017), benefits optimizers like Adam (Kingma and Ba, 2014). Similarly, Krizhevsky et al. (2012) showed that weight decay, when combined with Dropout (Hinton et al., 2012; Srivastava et al., 2014), reduces overfitting effectively. In addition, stochastic depth is another regularization technique designed to address challenges like vanishing gradients, diminishing feature reuse, and long training times in deep networks (Huang et al., 2016). Recent work (Zhou et al., 2025) demonstrated that stochastic depth is also effective in reducing information redundancy in ECG signals, improving performance on various ECG classification tasks. We explore these strategies in our study.
Following the general definition of regularization, initialization strategies can be considered a form of regularization. The pre-training paradigm, which has been successful in many ECG tasks (Zhou et al., 2025; Li et al., 2025; Coppola et al., 2024; McKeen et al., 2024), retains the pre-trained weights in the early layers of the model. However, the final classification head is typically initialized randomly. In this paper, we use linear probing to initialize this layer. Furthermore, early stopping can also be regarded as a regularization technique, and it is often used alongside other regularization methods (Goodfellow et al., 2016). We incorporate early stopping in our experiments as well.
3 Methodology
In this study, we propose a post-training strategy designed to enhance the classification performance of ECGFounder, a recently released state-of-the-art ECG foundation model (Li et al., 2025). The overall framework of the proposed approach is illustrated in Figure 2. Our framework builds upon the publicly available network architecture and pre-trained weights of ECGFounder, and introduces two key stages for downstream classification: an initialization stage and a regularization stage. The design and implementation of each stage are described in detail in the following subsections.
Figure 2: Overview of the proposed post-training framework for ECG classification.
3.1 Overview of ECGFounder
This subsection provides a concise overview of ECGFounder. The model is built upon the CNN-based Net1D architecture (Hong et al., 2020), adapted from RegNet (Radosavovic et al., 2020) and specifically optimized for ECG signal analysis. Its design incorporates stagewise scaling, bottleneck blocks (He et al., 2016), and channel-wise attention (Hu et al., 2018), enabling the model to capture both temporal dynamics and inter-lead spatial dependencies.
For pre-training, ECGFounder was built on part of the Harvard-Emory ECG Database (Koscova et al., 2024), which contains over 10 million ECGs across 150 diagnostic categories, using a multilabel setting that reflects the complexity of real-world clinical practice. To address incomplete annotations, a positive–unlabeled (PU) learning strategy (Zhao et al., 2022) was integrated into the loss function, allowing the model to account for unlabeled but potentially positive samples.
For downstream tasks, ECGFounder supports both linear probing and full fine-tuning, as proposed in Li et al. (2025). These training strategies are optimized to ensure robust performance across diverse ECG applications. Specifically, training is guided by a ReduceLROnPlateau learning rate scheduler with the AdamW optimizer. Model selection is based on validation macro Area Under the Receiver Operating Characteristic Curve (AUROC), aiming to enhance adaptability and reliability.
3.2 Post-training Strategy
As illustrated in Figure 2, our method builds upon ECGFounder and introduces a two-stage post-training strategy. The strategy is designed to enhance classification performance while requiring minimal modifications to the original model architecture. Specifically, it consists of an initialization stage and a regularization stage, which are described in detail below.
3.2.1 Initialization Stage
The initialization stage follows a standard linear probing approach. Specifically, all pretrained backbone parameters are frozen, and the linear classification head is replaced with a task-specific counterpart trained from scratch. This design allows the final layer to be rapidly aligned with the pretrained representations while providing a well-conditioned initialization for downstream classification. As discussed in Section 1, an appropriate initialization can facilitate task-specific adaptation and ultimately improve model performance. Following the general definition of regularization presented in Section 2, this step can also be viewed as a form of regularization.
3.2.2 Regularization Stage
The regularization stage involves full fine-tuning of the entire model following initialization, with multiple regularization techniques incorporated to improve generalization. The core component is the stochastic depth strategy (Huang et al., 2016), which has been shown to effectively reduce redundancy in ECG signals (Zhou et al., 2025). Specifically, stochastic depth is applied to all residual connections in ECGFounder. With a predefined drop-path rate $\lambda$, a binary random mask $r$ is defined as
$r=\mathbb{I}(U<1-\lambda),\quad U\sim\mathcal{U}(0,1),$ (1)
where $\mathcal{U}$ denotes the uniform distribution and $\mathbb{I}(\cdot)$ is the indicator function. Let $\mathbf{X}$ represent the input tensor, $F(\cdot)$ denote the block transformation, and $\widetilde{\mathbf{X}}$ the block output. The stochastic depth formulation during training is expressed as
$\widetilde{\mathbf{X}}=\frac{r}{1-\lambda}\cdot F(\mathbf{X})+\mathbf{X},$ (2)
where $r$ determines whether the transformation $F(\cdot)$ is active $(r=1)$ or skipped $(r=0)$, and the scaling factor $1/(1-\lambda)$ ensures that the expected output is preserved. During inference, the deterministic residual form is used:
$\widetilde{\mathbf{X}}=F(\mathbf{X})+\mathbf{X}.$ (3)
To further stabilize training, Dropout (Hinton et al., 2012; Srivastava et al., 2014) is applied to the activated outputs within all bottleneck blocks. In addition, a cosine annealing learning rate schedule (Loshchilov and Hutter, 2017) and weight decay (Hanson and Pratt, 1988; Loshchilov and Hutter, 2019) are employed to provide smooth decay and enhance robustness during optimization.
3.3 Ease of Implementation
To preserve the simplicity and accessibility of ECGFounder, we deliberately minimize modifications to the original framework. All proposed strategies avoid deep architectural changes and can be readily integrated into the existing implementation. In the initialization stage, our method can be directly applied within the open-source ECGFounder code. In the regularization stage, Dropout, weight decay the cosine annealing learning-rate schedule are standard functions in PyTorch. The stochastic depth operation is also straightforward to implement, requiring only a single-line modification within the bottleneck block:
  • out += identity
is replaced with
  • out = self.drop_path(out) + identity
Here, out and identity are tensors already defined in the implementation of Li et al. (2025), while self.drop_path() represents the stochastic depth strategy employed in this work. Overall, the proposed post-training strategy remains simple yet effective, preserving the ease of use of ECGFounder while providing enhanced performance.
4 Experiments
In this section, we present a comprehensive set of experiments conducted on the PTB-XL benchmark dataset (Strodthoff et al., 2021). We first describe the datasets and the experimental settings. We then evaluate the effectiveness of the proposed post-training strategy by comparing its performance against the suggested fine-tuning strategy (Li et al., 2025) and several state-of-the-art methods, including recent ECG foundation models and task-specific approaches that have been carefully optimized. Additionally, we conduct experiments to further investigate the properties of the proposed method, including performance comparisons under varying training epochs and sample sizes, as well as an ablation study.
4.1 Datasets
The PTB-XL database is a widely used open-source dataset for evaluating ECG model performance, notable for its relatively large sample size, high-quality annotations and clearly defined usage guidelines (Wagner et al., 2020; Strodthoff, 2024). It contains 21,837 clinical 12-lead ECG recordings, each 10 seconds in duration, collected from 18,885 patients across a broad age range (0–95 years). Each recording was annotated by up to two cardiologists, with labels drawn from a set of 71 distinct ECG statements. These statements can be organized into 12 rhythm labels and 44 diagnostic labels, which can further be hierarchically grouped into 24 subclasses. Multiple labels may be associated with a single recording, reflecting the inherently multi-label nature of clinical ECG interpretation. Based on this labeling system, we consider four multi-label classification tasks: (1) all-71, which uses the complete set of 71 labels; (2) rhythm-12, which focuses on 12 rhythm categories; (3) diagnostic-44, which focuses on 44 diagnostic categories; and (4) subclass-23, which focuses on 23 hierarchical subclasses derived from the diagnostic labels. As discussed in Section 2, PTB-XL provides predefined splits into training, validation, and testing sets (approximately 8:1:1), enabling consistent and fair comparisons across methods (Strodthoff et al., 2021; Strodthoff, 2024). We follow this standard data splitting in all experiments.
4.2 Experimental Settings
In the initialization stage, we follow the training recipe recommended in Li et al. (2025), except for a shorter training length, since the largest validation macro Area Under the Receiver Operating Characteristic Curve (AUROC) is typically achieved in the early epochs. Specifically, the model is trained for 30 epochs with a batch size of 256, using AdamW with a weight decay of 0.1 and an initial learning rate of $5\times 10^{-3}$. A ReduceLROnPlateau scheduler (maximization mode) monitors macro AUROC on the validation set and reduces the learning rate by a factor of 0.1 if no improvement is observed for 10 consecutive epochs.
In the regularization stage, we also train for 30 epochs with a batch size of 256. Optimization is performed using AdamW with a weight decay of 0.05 and a learning rate of $5\times 10^{-4}$. A CosineAnnealingLR scheduler is adopted with a minimum learning rate of $10^{-5}$. In addition, we apply a droppath rate of 0.5 and a dropout rate of 0.05.
For both stages, the model is evaluated on the validation set at the end of each epoch, and the checkpoint with the highest macro AUROC is selected for final evaluation. This strategy serves as a form of early stopping regularization.
4.3 Comparison with Other Methods
This section presents a comprehensive comparison between the proposed post-training method and a range of baseline approaches, including the suggested baseline fine-tuning strategy, referred to ECGFounder (Li et al., 2025). To distinguish the post-training variant, we denote our method as ECGFounder-PT (Post-Training). Additional baselines include the foundation model HuBERT-ECG-BASE (Coppola et al., 2024) and task-specific and advanced models, such as inception1d (Ismail Fawaz et al., 2020), resnet1d101 (Strodthoff et al., 2021), resnet1d_wang (Wang et al., 2017), ensemble-7-models (Strodthoff et al., 2021), SPACETIME (Zhang et al., 2023), MSW-Transformer (Cheng et al., 2023), MULTIRESNET (Shi et al., 2023), and Chimera (Behrouz et al., 2024). The reported performance values are derived from (Strodthoff et al., 2021; Zhang et al., 2023; Cheng et al., 2023; Shi et al., 2023; Coppola et al., 2024; Behrouz et al., 2024).
Method all-71 diagnostic-44 subclass-23 rhythm-12
inception1d 0.925 0.931 0.930 0.953
xresnet1d101 0.925 0.937 0.929 0.957
resnet1d_wang 0.919 0.936 0.928 0.946
ensemble-7-models 0.929 0.939 0.933 0.965
MULTIRESNET 0.938 0.939 0.934 0.975
MSW-Transformer - 0.924 0.927 0.925
SPACETIME 0.936 0.941 0.933 0.967
Chimera 0.941 0.947 0.935 0.975
HuBERT-ECG-BASE 0.903 0.917 0.917 0.953
ECGFounder
0.915
(0.904, 0.924)
0.916
(0.902, 0.928)
0.927
(0.913, 0.939)
0.964
(0.948, 0.977)
ECGFounder-PT (ours)
0.944
(0.936, 0.951)
0.946
(0.939, 0.951)
0.938
(0.927, 0.947)
0.979
(0.965, 0.988)
Table 1: Performance comparison in terms of macro AUROC. Best and second-best results are highlighted in bold and underlined, respectively.
Table 1 presents the macro AUROC of all methods, along with 95% confidence intervals computed using 1000 bootstrapped samples. ECGFounder-PT achieves the highest AUROC on three tasks: all-71 (0.944), subclass-23 (0.938), and rhythm-12 (0.979), and the second-highest on diagnostic-44 (0.946). Compared with the original ECGFounder and HuBERT-ECG-BASE, ECGFounder-PT improves performance by 1.2%–3.3% and 2.3%-4.5%, respectively, demonstrating the effectiveness of the post-training adaptation.
In addition to foundation models, ECGFounder-PT also outperforms several recent state-of-the-art task-specific approaches. Compared to CNN-based baselines (inception1d, resnet1d101, resnet1d_wang, and ensemble-7-models) from Strodthoff et al. (2021), ECGFounder-PT yields macro AUROC gains of 0.7%–3.5%.When compared to multi-scale architectures like MULTIRESNET and MSW-Transformer, which explicitly exploit the hierarchical properties of ECG signals, it improves performance by 0.4%–5.8%. Even compared with recent state space models (SPACETIME and Chimera), ECGFounder-PT achieves consistent gains of 0.3%–1.2%, except for diagnostic-44, where Chimera slightly outperforms it by 0.1%. Collectively, these results highlight the robustness of ECGFounder-PT and its ability to bridge the adaptation gap of foundation models, achieving balanced generalization across diverse ECG classification tasks.
Although AUROC is a standard benchmark metric, it may be less sensitive in the presence of class imbalance. To address this, we additionally report macro Area Under the Precision-Recall Curve (AUPRC) in Table 2, with 95% confidence intervals computed using 1000 bootstrapped samples. Since AUPRC results for other methods are not available, the comparison is limited to ECGFounder and ECGFounder-PT. Across all tasks, ECGFounder-PT consistently outperforms ECGFounder, with relative gains ranging from 5.3% to 20.9%, underscoring the effectiveness of the post-training strategy, particularly in imbalanced data scenarios.
Method all-71 diagnostic-44 subclass-23 rhythm-12
ECGFounder
0.356
(0.349, 0.397)
0.321
(0.310, 0.358)
0.486
(0.464, 0.533)
0.604
(0.547, 0.687)
ECGFounder-PT (ours)
0.407
(0.397, 0.447)
0.388
(0.374, 0.438)
0.512
(0.480, 0.561)
0.645
(0.593, 0.739)
Table 2: Performance comparison in terms of macro AUPRC. Bold numbers indicate the best performance for each task.
4.4 Performance with Varying Training Epochs
We report the performance of the proposed post-training method (ECGFounder-PT) and the baseline fine-tuning strategy (ECGFounder) across epochs on the PTB-XL all-71 task in Figure 3. Since the validation and testing sets show similar trends, we present the results from the testing set for clarity. This analysis underscores the robustness and generalization of ECGFounder-PT compared to ECGFounder.
As shown in Figure 3(b), ECGFounder-PT consistently outperforms ECGFounder across both macro AUROC and macro AUPRC. The advantage is evident from the first epoch: ECGFounder-PT achieves a macro AUROC of 0.922 and a macro AUPRC of 0.384, compared to the baseline’s 0.762 and 0.118, respectively. This immediate superiority demonstrates the effectiveness of the post-training strategy in rapidly adapting the pre-trained model to downstream diagnostic tasks, leveraging prior knowledge more efficiently than the baseline.
The most notable difference lies in convergence dynamics and final performance. The baseline strategy exhibits a prolonged and unstable learning curve: its macro AUROC on the testing set rises gradually, peaking at 0.919 at epoch 16, but fails to stabilize, fluctuating between 0.901 and 0.919 from epochs 16 to 30. Similar oscillations are observed on the validation set, suggesting overfitting and limited generalization. In contrast, the proposed post-training strategy converges rapidly and smoothly to a higher, stable plateau. Macro AUROC surpasses 0.940 by epoch 8 and remains around 0.944 thereafter, with negligible variance. This stable convergence indicates that ECGFounder-PT effectively regularizes the optimization process, guiding the model toward a broad and generalizable optimum.
As shown in Figure 3(d), a similar trend is observed for the more challenging macro AUPRC, which is more informative for assessing performance on imbalanced datasets. The baseline strategy exhibits slow and unsteady improvement, reaching a maximum test macro AUPRC of 0.368. In contrast, ECGFounder-PT achieves a rapid and substantial gain, stabilizing around 0.435. This 18.2% relative improvement is particularly clinically meaningful, reflecting a markedly enhanced ability to identify true positives in imbalanced datasets, a common scenario in real-world ECG diagnosis.
Figure 3: Comparison of AUROC and AUPRC between the baseline training recipe and the proposed post-training strategy across epochs on PTB-XL validation and testing sets.
4.5 Performance with Varying Training Sample Sizes
Figure 4 compares the proposed post-training method (ECGFounder-PT) with the baseline fine-tuning strategy (ECGFounder) on the PTB-XL all-71 task, using varying training sample sizes from 10% to 100% of the available data. Macro AUROC and macro AUPRC are used as evaluation metrics. As in Section 4.4, since validation and test sets exhibit similar patterns, we present results from the testing set for clarity. This analysis explores the applicability of the proposed method in real-world scenarios where the availability of training data varies across medical institutions.
Overall, ECGFounder-PT consistently outperforms ECGFounder across all sample sizes, with improvements ranging from 3.0% to 9.1% in macro AUROC and 10.7% to 34.9% in macro AUPRC (Figures 4(b) and 4(d)).
The largest performance gains occur under limited data conditions. With only 10% of the training set, ECGFounder-PT achieves a macro AUROC of 0.899 compared to 0.824 for ECGFounder (+9.1%) and a macro AUPRC of 0.294 compared to 0.218 (+34.9%). These results demonstrate the substantial advantage of the post-training strategy in data-scarce environments, a common challenge in clinical practice.
As training data increases, the relative improvements decrease but remain consistent. For instance, at 30% of the training data, ECGFounder-PT achieves an AUROC of 0.925 and an AUPRC of 0.364, outperforming ECGFounder’s 0.877 and 0.301 (+5.5% and +21.0%, respectively). Even at higher data availability (e.g., 90%), ECGFounder-PT maintains an edge, reaching 0.942 AUROC and 0.402 AUPRC compared to 0.915 and 0.363, corresponding to relative gains of 2.9% and 10.7%.
These findings highlight the robustness of ECGFounder-PT: the post-training strategy yields consistent performance gains across diverse data regimes, with the most pronounced benefits under data scarcity. This property underscores its practical value for real-world deployment across diverse healthcare environments.
Figure 4: Comparison of AUROC and AUPRC between the baseline training recipe and the proposed post-training strategy across varying training sample sizes on PTB-XL validation and testing sets.
4.6 Ablation Study
To assess the contribution of each component in the proposed post-training strategy, we conducted an ablation study on the PTB-XL all-71 task. The results are reported in Table 3. In this analysis, “random initialization” refers to training the model from scratch with default weight initialization. For all other cases, the initialization was derived from ECGFounder, which was pre-trained on more than 7 million ECG recordings. “without droppath” indicates the removal of the stochastic depth operation, while “without linear probing” excludes the preview linear probing initialization. Likewise, “without dropout” and “without cosine schedule” denote the exclusion of dropout regularization and cosine annealing learning rate decay, respectively.
As shown in Table 3, every component in the proposed strategy contributes to the overall performance improvement. Among them, the stochastic depth operation and the preview linear probing initialization are the most important. Removing either of these components results in substantial performance degradation: the AUPRC decreases by approximately 4.1%–5.1% on the validation set and 0.7%–4.4% on the testing set, while AUROC decreases by 0.3%–0.4% and 1.1%–1.6% on the validation and testing sets, respectively. In contrast, the absence of dropout or cosine learning rate schedule leads to only marginal differences, suggesting that these techniques primarily provide auxiliary benefits.
It is also noteworthy that, even under random initialization, the proposed post-training strategy achieves performance comparable to the baseline strategy that leverages pre-training on more than 7 million ECGs. This demonstrates the robustness of the proposed approach and underscores its potential value for advancing ECG foundation models.
Table 3: Ablation study on the PTB-XL dataset (all-71 task). The table reports the effect of removing individual components from the proposed post-training strategy. Best and second-best results are highlighted in bold and underlined, respectively.
Method Setting Validation set Testing set
AUROC AUPRC AUROC AUPRC
Baseline Strategy random initialization 0.904 0.312 0.902 0.317
pre-training 0.924 0.348 0.915 0.356
Proposed Strategy random initialization 0.925 0.359 0.919 0.349
without droppath 0.940 0.417 0.929 0.389
without linear probing 0.939 0.413 0.934 0.404
without dropout 0.943 0.428 0.943 0.405
without cosine schedule 0.943 0.425 0.942 0.407
all components 0.943 0.435 0.944 0.407
5 Discussion
1. Performance Gains on ECG Foundation Models
In this study, we evaluated the effectiveness of the proposed post-training strategy on four PTB-XL multi-label ECG classification benchmark tasks. The proposed method demonstrates consistent improvements over recent open-source ECG foundation models, confirming its ability to enhance generalization across diverse ECG classification tasks. The macro AUROC values of 0.944 (all-71), 0.946 (diagnostic-44), 0.938 (subclass-23), and 0.979 (rhythm-12) indicate 1.2%–3.3% gains over the baseline fine-tuning strategy (Li et al., 2025) and 2.3%-4.5% over another foundation model HuBERT-ECG-BASE (Coppola et al., 2024). Improvements in the macro AUPRC, ranging from 5.3% to 20.9%, compared to the baseline fine-tuning strategy, further highlight the model’s robustness in the context of imbalanced clinical datasets.
2. Comparison with State-of-the-Art
It is worth noting that the proposed method consistently outperforms task-specific and advanced architectures. Baseline methods such as inception1d, xresnet1d101, resnet1d_wang, and ensemble-7-models (Strodthoff et al., 2021) exhibit competitive or even superior performance to recent ECG foundation models, likely due to their specialized design for ECG signals. Similarly, the multi-scale architecture MULTIRESNET (Shi et al., 2023) performs well, reflecting the inherent multi-scale structure in ECG data, while advanced models like SPACETIME (Zhang et al., 2023) and Chimera (Behrouz et al., 2024) highlight the significance of carefully designed architectures for ECG time-series data. These task-specific and advanced models reveal a deeper understanding of ECG signals, as they are tailored for the specific challenges of ECG classification. In contrast, foundation models like ECGFounder and HuBERT-ECG-BASE prioritize broad adaptability, which may not rely on a strong inductive bias specific to some ECG tasks. This flexibility allows foundation models to be applied across a wide range of tasks, but in certain cases, it can result in lower performance compared to specialized models, especially in the absence of an appropriate post-training strategy. As demonstrated in both our study and the work of Coppola et al. (2024), foundation models often lag behind task-specific and advanced architectures, even when pre-trained on large ECG datasets. This performance gap underscores the critical need for a robust post-training strategy. Our results, shown in Table 1, clearly illustrate that incorporating an effective post-training approach significantly enhances the performance of ECG foundation models, enabling them to match or even surpass task-specific and advanced architectures.
3. Sample Efficiency and Limited Data Scenarios
The proposed post-training strategy demonstrates substantial benefits across a wide range of sample sizes, with particularly significant improvements in scenarios where data is limited. As shown in Figure 4, when trained with only 10% of the dataset, approximately 1,700 ECG samples, the proposed method achieves a macro AUROC of 0.899 and a macro AUPRC of 0.294, compared to 0.824 and 0.218 for the baseline strategy (Li et al., 2025), resulting in relative improvements of 9.1% and 34.9%, respectively. Even with 90% of the training data, the post-training strategy continues to deliver consistent gains, with macro AUROC and AUPRC improvements of 2.9% and 10.7%, respectively. These findings highlight the effectiveness of the post-training approach in reducing the dependency on large annotated datasets, which is particularly beneficial in clinical ECG applications where labeled data is often scarce. This ability to enhance model performance with limited data underscores the practical utility of the proposed strategy in real-world medical settings, where access to extensive labeled ECG datasets may be constrained.
4. Training Stability and Convergence
A key factor driving the observed improvements is the enhanced training stability and convergence behavior facilitated by the proposed post-training strategy. Specifically, the method converges rapidly, surpassing a macro AUROC of 0.940 by epoch 8 and stabilizing at a plateau around 0.944 for the remainder of the training. In contrast, the baseline strategy exhibits slower convergence, with AUROC oscillating between 0.901 and 0.919 from epochs 16 to 30. Similarly, macro AUPRC reaches a stable value of 0.435, representing an 18.2% relative improvement over the baseline. These findings indicate that the post-training strategy effectively regularizes the optimization process, steering the model towards a more generalizable solution while mitigating overfitting. This enhanced stability not only accelerates convergence but also contributes to the overall robustness of the model, ensuring better performance across varying training conditions.
5. Ablation Study Insights
The ablation study highlights the contribution of individual components to overall performance. Stochastic depth and preview linear probing emerged as the most influential factors, with their removal leading to a decrease in macro AUPRC by 4.1%–5.1% on the validation set and 0.7%–4.4% on the testing set. The importance of stochastic depth has also been confirmed in a recent study on ECG classification (Zhou et al., 2025), underscoring its role in enhancing generalization. Preview linear probing, on the other hand, provides better initialization for the classification head, thereby boosting performance as expected. Conceptually, this strategy resonates with initialization approaches in statistical learning. For example, motivated by Zhou et al. (2013), Zhou et al. (2024) propose a sequential downsizing initialization strategy for tensor regression, which essentially performs lightweight training to induce a bias aligned with the target data. Although theoretical understanding of deep learning still lags behind its empirical success, insights from traditional statistical models may inspire more principled approaches to enhance modern deep learning methods.
6. Evaluation Metrics
Another important observation pertains to the performance metrics used. While macro AUROC remains the standard metric in ECG benchmarks (Strodthoff et al., 2021), it is relatively insensitive to class imbalance, which is a common characteristic of clinical ECG datasets. In contrast, macro AUPRC is more sensitive to this imbalance, as it captures the precision-recall trade-off (Ozenne et al., 2015). This distinction is evident in most of our experiments. For instance, in the four PTB-XL multi-label tasks examined in this study, the proposed method achieves improvements of 1.2%–3.3% in macro AUROC compared to the baseline, whereas the improvement in macro AUPRC ranges from 5.3% to 20.9%. On the one hand, the substantial gains in macro AUPRC highlight the efficacy of the post-training strategy. On the other hand, the notable discrepancy between macro AUROC and AUPRC underscores an important insight: even a small difference in macro AUROC can correspond to significant variations in macro AUPRC, indicating that different methods may exhibit considerable differences in their ability to detect true positives in the presence of a large number of negatives.
7. Implications and Future Directions
Our findings have important implications for the design and clinical deployment of ECG foundation models. While recent successes of foundation models in natural language processing have inspired similar approaches for health data (Clifford, 2024), ECG foundation models face unique challenges due to the domain-specific characteristics of ECG signals, such as high information redundancy and physiological variability (Zhou et al., 2025). Recent studies have reported that existing ECG foundation models achieve sub-optimal performance on the PTB-XL benchmark (Coppola et al., 2024), raising concerns regarding their direct clinical applicability. Our results indicate that this limitation may be attributed to the absence of an effective post-training strategy. By addressing this gap with our proposed post-training strategy, the ECG foundation model can achieve strong performance on downstream tasks. This highlights the critical role of post-training in enabling real-world clinical adoption of foundation models.
6 Conclusion
In this study, we propose a simple yet effective post-training strategy for the recently released ECG foundation model, ECGFounder. Experiments on the PTB-XL benchmark demonstrate that the proposed method outperforms both the baseline fine-tuning strategy and several recent state-of-the-art models. Key components, including stochastic depth and preview linear probing, were identified as critical drivers of performance improvement. Additionally, the proposed method exhibits greater sample efficiency and training stability. Our findings underscore the potential of post-training strategies to bridge the adaptation gap for ECG foundation models, highlighting a promising direction to further advance their development.